# Sampling as Bandits: Evaluation-Efficient Design for Black-Box Densities

### v1

@@ B.2 Proof of Theorem˜1
$\left|\int_{\Theta}g(\theta)d\theta-\frac{\operatorname{Vol}(\Theta)}{K}\sum_{n=1}^{K}g(\theta_{n})\right|=\operatorname{Vol}(\Theta)\underbrace{\left|\int_{[0,1]^{d}}g\circ\tau(\eta)d\eta-\frac{1}{K}\sum_{n=1}^{K}g\circ\tau(\eta_{n})\right|}_{=:(*)}$ (60)
where we used that $\prod_{i=1}^{d}|(\partial/\partial\eta_{(i)})\tau_{i}(\eta_{(i)})|=\prod_{i=1}^{d}|b_{i}-a_{i}|=\operatorname{Vol}(\Theta)$ for the change of variables. Recall the star discrepancy $\operatorname{D}^{\dagger}$ for the unit cube $0,1]^{d}$ defined in (25) in Section˜A.2. It follows from the Koksma-Hlawka inequality (Dick and Pillichshammer, 2010, Proposition 2.18) that
$(*)\leq V^{\dagger}(g\circ\tau)\operatorname{D}^{\dagger}(\{\eta_{n}\}_{n=1}^{K}).$ (61)
@@
$V^{\dagger}(g\circ\tau):=\sum_{\alpha\neq 0}\int_{[0,1]^{\alpha}}\left|\partial^{\alpha}g\circ\tau(\eta_{1}^{\alpha})\right|d\eta^{\alpha}.$ (62)
@@
$\partial^{\alpha}g\circ\tau(\eta_{1}^{\alpha})=\partial^{\alpha}g(\theta)\big{|}_{\theta=\tau(\eta_{1}^{\alpha})}\left(\prod_{i\in I_{\alpha}}\partial_{i}\tau_{i}(\eta_{(i)})\right)=\partial^{\alpha}g(\theta)\big{|}_{\theta=\tau(\eta_{1}^{\alpha})}\cdot\operatorname{Vol}(\Theta^{\alpha})$ (63)
where we used $\prod_{i\in I_{\alpha}}\partial_{i}\tau_{i}(\eta_{(i)})=\prod_{i\in I_{\alpha}}b_{i}-a_{i}=\operatorname{Vol}(\Theta^{\alpha})$. Denote by $h(\theta)$ the mixed partial derivative $\partial^{\alpha}g(\theta)$ to see that $h(\tau(\eta_{1}^{\alpha}))=\partial^{\alpha}g(\theta)|_{\theta=\tau(\eta_{1}^{\alpha})}$. Combining (62) and (63) gives that
$V^{\dagger}(g\circ\tau)=\sum_{\alpha\neq 0}\int_{[0,1]^{\alpha}}\operatorname{Vol}(\Theta^{\alpha})\left|h(\tau(\eta_{1}^{\alpha}))\right|d\eta^{\alpha}.$ (64)
@@
$V^{\dagger}(g\circ\tau)=\sum_{\alpha\neq 0}\int_{\Theta^{\alpha}}\operatorname{Vol}(\Theta^{\alpha})\left|h(\tau(\eta_{1}^{\alpha}))\right|d\eta^{\alpha}=\sum_{\alpha\neq 0}\int_{\Theta^{\alpha}}\left|\partial^{\alpha}g(\theta_{*}^{\alpha})\right|d\theta^{\alpha}=V(g).$ (65)
where $\operatorname{Vol}(\Theta^{\alpha})$ is cancelled by the change of variables. Therefore, the inequality (61) is rewritten as
$(*)\leq V^{*}(g)\operatorname{D}^{\dagger}(\{\eta_{n}\}_{n=1}^{K})$ (66)
@@
$(*_{1})$ $=\sup_{\|f\|_{\mathcal{H}_{\kappa}\leq 1}}\left|\int_{\Theta}f(\theta)q(\theta)d\theta-\frac{\operatorname{Vol}(\Theta)}{N+M}\sum_{n=1}^{N+M}f(\theta_{n})q(\theta_{n})\right|\leq\operatorname{Vol}(\Theta)\underbrace{\sup_{\|f\|_{\mathcal{H}_{\kappa}\leq 1}}V(f\cdot q)}_{=:(*_{11})}\operatorname{D}^{*}(\{\theta_{n}\}_{n=1}^{M+N})$ (68)
@@
$V(f\cdot q)$ $\leq\sum_{\alpha\neq 0}\operatorname{Vol}(\Theta^{\alpha})\sup_{\theta\in\Theta^{\alpha}}\left|\partial^{\alpha}(f(\theta_{*}^{\alpha})q(\theta_{*}^{\alpha}))\right|\leq\sum_{\alpha\neq 0}\operatorname{Vol}(\Theta^{\alpha})\underbrace{\sup_{\theta\in\Theta}\left|\partial^{\alpha}(f(\theta)q(\theta))\right|}_{=:(*_{12})}.$ (69)
@@
$\partial^{\alpha}(f(\theta)q(\theta))=\sum_{\beta\leq\alpha}(\partial^{\alpha-\beta}f(\theta))(\partial^{\beta}q(\theta))$ (70)
@@
$(*_{12})\leq\sup_{\theta\in\Theta}\sum_{\beta\leq\alpha}\left|(\partial^{\alpha-\beta}f(\theta))(\partial^{\beta}q(\theta))\right|\leq\sum_{\beta\leq\alpha}C_{q}^{\beta}.$ (71)
@@
$(*_{11})\leq\sum_{\alpha\neq 0}\operatorname{Vol}(\Theta^{\alpha})\sum_{\beta\leq\alpha}C_{q}^{\beta}=:A_{1}$ (72)
@@
$(*_{2})=\left|\int_{\Theta}q(\theta)d\theta-\frac{\operatorname{Vol}(\Theta)}{N+M}\sum_{n=1}^{N+M}q(\theta_{n})\right|\leq\operatorname{Vol}(\Theta)V(q)\operatorname{D}^{*}(\{\theta_{n}\}_{n=1}^{M+N})$ (73)
@@
$V(q)$ $\leq\sum_{\alpha\neq 0}\operatorname{Vol}(\Theta^{\alpha})\sup_{\theta\in\Theta^{\alpha}}\left|\partial^{\alpha}q(\theta_{*}^{\alpha})\right|\leq\sum_{\alpha\neq 0}\operatorname{Vol}(\Theta^{\alpha})\sup_{\theta\in\Theta}\left|\partial^{\alpha}q(\theta)\right|.$ (74)
@@
$V(q)$ $\leq\sum_{\alpha\neq 0}\operatorname{Vol}(\Theta^{\alpha})C_{q}^{\alpha}=:A_{2}$ (75)
@@
$(*)$ $\leq\operatorname{Vol}(\Theta)A_{1}B_{1}\operatorname{D}^{*}(\{\theta_{n}\}_{n=1}^{M+N})+\operatorname{Vol}(\Theta)A_{2}B_{2}\operatorname{D}^{*}(\{\theta_{n}\}_{n=1}^{M+N})+B_{3}\frac{M}{N+M}.$ (76)
@@ B.3 Proof of Corollary˜1
Recall that $\operatorname{D}^{\dagger}$ is the star discrepancy for the unit cube $0,1]^{d}$ defined in (25) in Section˜A.2. The scaled Halton sequence $\{\theta_{i}\}_{i=1}^{\infty}$ is defined as an image of the original Halton sequence $\{\eta_{i}\}_{i=1}^{\infty}$ under the bijective affine map. By Lemma˜3, we have $\operatorname{D}^{*}\left(\{\theta_{n}\}_{n=1}^{K}\right)=\operatorname{D}^{\dagger}\left(\{\eta_{n}\}_{n=1}^{K}\right)$ for any number of points $K$. By Theorem 3.6 of Niederreiter (1992), there exists some constant $C>0$ s.t. the star discrepancy $\operatorname{D}^{\dagger}$ of the original Halton sequence $\{\eta_{n}\}_{n=1}^{K}$ satisfies that
$\operatorname{D}^{\dagger}\left(\{\eta_{n}\}_{n=1}^{K}\right)\leq C\frac{\log(K)^{d}}{K}.$ (77)
@@ B.4 Proof of Corollary˜2
By assumption, the kernel $\kappa$ is bounded, continuous, and integrally strictly positive definite. Hence, the resulting MMD metrises the weak convergence by Theorem 7 of Simon-Gabriel et al. (2023). By Corollary˜1, the MMD between $\delta_{N}$ and $p$ converges to zero, concluding the proof. ∎
B.5 Proof of Theorem˜2
The proof uses Lemma˜4 presented in Section˜B.2, which holds for any proposal sequence.
@@ B.5 Proof of Theorem˜2
$I_{M+N}(f)$ $=\left|\int_{\Theta}f(\theta)g(\theta)u(\theta)d\theta-\frac{1}{N+M}\sum_{n=1}^{N+M}f(\theta_{n})g(\theta_{n})\right|.$ (79)
@@
$(*_{1})$ $=\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}\left|\mathbb{E}_{\theta\sim u}\left[f(\theta)g(\theta)\right]-\mathbb{E}_{\theta\sim u_{M+N}}\left[f(\theta)g(\theta)\right]\right|=\sup_{h\in\mathcal{H}}\left|\mathbb{E}_{\theta\sim u}\left[h(\theta)\right]-\mathbb{E}_{\theta\sim u_{M+N}}\left[h(\theta)\right]\right|.$ (80)
@@
$\mathcal{R}_{M+N}(\mathcal{H})$ $:=\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}}\mathbb{E}_{\epsilon_{1},\dots,\epsilon_{M+N}}\left[\sup_{h\in\mathcal{H}}\left|\frac{1}{M+N}\sum_{n=1}^{M+N}\epsilon_{n}h(\theta_{n})\right|\right]$ (81)
$=\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}}\mathbb{E}_{\epsilon_{1},\dots,\epsilon_{M+N}}\left[\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}\left|\frac{1}{M+N}\sum_{n=1}^{M+N}\epsilon_{n}f(\theta_{n})g(\theta_{n})\right|\right].$ (82)
@@
$\sup_{\theta\in\Theta}|h(\theta)|\leq\sup_{\theta\in\Theta}|f(\theta)|\sup_{\theta\in\Theta}|g(\theta)|\leq\sup_{\theta\in\Theta}|g(\theta)|=b.$ (83)
It then follows from Theorem 4.10 of Wainwright (2019) that
$\mathbb{P}\left((*_{1})\leq 2\mathcal{R}_{M+N}(\mathcal{H})+\delta\right)\geq 1-\exp\left(-\frac{(M+N)\delta^{2}}{2b^{2}}\right)$ (84)
@@
$\mathcal{R}_{M+N}(\mathcal{H})$ $=\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}}\mathbb{E}_{\epsilon_{1},\dots,\epsilon_{M+N}}\left[\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}\left|\frac{1}{M+N}\sum_{n=1}^{M+N}\epsilon_{n}\langle f(\cdot),\kappa(\cdot,\theta_{n})\rangle_{\mathcal{H}_{\kappa}}g(\theta_{n})\right|\right]$ (85)
$=\frac{1}{M+N}\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}}\mathbb{E}_{\epsilon_{1},\dots,\epsilon_{M+N}}\left[\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}\left|\left\langle f(\cdot),\sum_{n=1}^{M+N}\epsilon_{n}\kappa(\cdot,\theta_{n})g(\theta_{n})\right\rangle_{\mathcal{H}_{\kappa}}\right|\right].$ (86)
@@
$\mathcal{R}_{M+N}(\mathcal{H})$ $=\frac{1}{M+N}\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}}\mathbb{E}_{\epsilon_{1},\dots,\epsilon_{M+N}}\left[\left\|\sum_{n=1}^{M+N}\epsilon_{n}\kappa(\cdot,\theta_{n})g(\theta_{n})\right\|_{\mathcal{H}_{\kappa}}\right]$ (87)
$\leq\frac{1}{M+N}\sqrt{\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}}\mathbb{E}_{\epsilon_{1},\dots,\epsilon_{M+N}}\left[\left\|\sum_{n=1}^{M+N}\epsilon_{n}\kappa(\cdot,\theta_{n})g(\theta_{n})\right\|_{\mathcal{H}_{\kappa}}^{2}\right]}$ (88)
$=\frac{1}{M+N}\sqrt{\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}}\mathbb{E}_{\epsilon_{1},\dots,\epsilon_{M+N}}\left[\sum_{n=1}^{M+N}\sum_{m=1}^{M+N}\epsilon_{n}\epsilon_{m}g(\theta_{n})\kappa(\theta_{n},\theta_{m})g(\theta_{m})\right]}.$ (89)
@@
$\mathcal{R}_{M+N}(\mathcal{H})$ $\leq\frac{1}{M+N}\sqrt{\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}}\left[\sum_{n=1}^{M+N}g(\theta_{n})\kappa(\theta_{n},\theta_{n})g(\theta_{n})\right]}$ (90)
$\leq\frac{1}{M+N}\sqrt{(M+N)b^{2}}\leq\frac{b}{\sqrt{M+N}}$ (91)
where the second inequality follows from that $\kappa$ is bounded by $1$ and that $g(\theta)\leq\sup_{\theta\in\Theta}|g(\theta)|=b$. Plugging this upper bound of $\mathcal{R}_{M+N}(\mathcal{H})$ in the concentration inequality (84) gives that
$\mathbb{P}\left((*_{1})\leq\frac{2b}{\sqrt{M+N}}+\delta\right)\geq 1-\exp\left(-\frac{(M+N)\delta^{2}}{2b^{2}}\right)\geq 1-2\exp\left(-\frac{(M+N)\delta^{2}}{2b^{2}}\right)$ (92)
@@
$\mathbb{P}\left((*_{1})\leq\frac{2b+\sqrt{2\log(2/\epsilon)}b}{\sqrt{M+N}}\right)\geq 1-\epsilon$ (93)
@@
$(*_{2})$ $=\left|\mathbb{E}_{\theta\sim u}\left[g(\theta)\right]-\mathbb{E}_{\theta\sim u_{M+N}}\left[g(\theta)\right]\right|=\left|\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}\overset{i.i.d.}{\sim}u}[F(\theta_{1},\dots,\theta_{M+N})]-F(\theta_{1},\dots,\theta_{M+N})\right|.$ (94)
Our aim is to apply the bounded difference inequality in Corollary 2.21 of Wainwright (2019). The function $F$ admits the ‘bounded difference’ with respect to each $n$-th argument $\theta_{n}$. To see this, fix all arguments $\theta_{1},\dots,\theta_{M+N}$, except the $n$-th argument $\theta_{n}$, and observe for any $n$ that
$\left|F(\theta_{1},\dots,\theta_{n},\dots,\theta_{M+N})-F(\theta_{1},\dots,\theta_{n}^{\prime},\dots,\theta_{M+N})\right|=\frac{1}{M+N}\left|g(\theta_{n})-g(\theta_{n}^{\prime})\right|\leq\frac{2b}{M+N}$ (95)
where $\theta_{n}$ and $\theta_{n}^{\prime}$ are arbitrary. Thus we can apply Corollary 2.21 of Wainwright (2019) to see that
$\mathbb{P}\left((*_{2})\leq\delta\right)\geq 1-2\exp\left(-\frac{(M+N)\delta^{2}}{2b^{2}}\right)$ (96)
@@
$\mathbb{P}\left((*_{2})\leq\frac{\sqrt{2\log(2/\epsilon)}b}{\sqrt{M+N}}\right)\geq 1-\epsilon.$ (97)
@@
$\mathbb{P}\left(d(p,\delta_{N})\leq B_{1}\frac{2b+\sqrt{2\log(2/\epsilon)}b}{\sqrt{M+N}}+B_{2}\frac{\sqrt{2\log(2/\epsilon)}b}{\sqrt{M+N}}+B_{3}\frac{M}{N+M}\right)\geq 1-\epsilon$ (98)

### v2

@@ B.1 Proof of Theorem˜1
$\left|\int_{\Theta}g(\theta)d\theta-\frac{\operatorname{Vol}(\Theta)}{K}\sum_{n=1}^{K}g(\theta_{n})\right|=\operatorname{Vol}(\Theta)\underbrace{\left|\int_{[0,1]^{d}}g\circ\tau(\eta)d\eta-\frac{1}{K}\sum_{n=1}^{K}g\circ\tau(\eta_{n})\right|}_{=:(*)}$ (S33)
where we used that $\prod_{i=1}^{d}|(\partial/\partial\eta_{(i)})\tau_{i}(\eta_{(i)})|=\prod_{i=1}^{d}|b_{i}-a_{i}|=\operatorname{Vol}(\Theta)$ for the change of variables. Recall the star discrepancy $\operatorname{D}^{\dagger}$ for the unit cube $0,1]^{d}$ defined in (S8) in Section˜A.2. It follows from the Koksma-Hlawka inequality (Dick2010, Proposition 2.18) that
$(*)\leq V^{\dagger}(g\circ\tau)\operatorname{D}^{\dagger}(\{\eta_{n}\}_{n=1}^{K}).$ (S34)
@@
$V^{\dagger}(g\circ\tau):=\sum_{\alpha\neq 0}\int_{[0,1]^{\alpha}}\left|\partial^{\alpha}g\circ\tau(\eta_{1}^{\alpha})\right|d\eta^{\alpha}.$ (S35)
@@
$\partial^{\alpha}g\circ\tau(\eta_{1}^{\alpha})=\partial^{\alpha}g(\theta)\big|_{\theta=\tau(\eta_{1}^{\alpha})}\left(\prod_{i\in I_{\alpha}}\partial_{i}\tau_{i}(\eta_{(i)})\right)=\partial^{\alpha}g(\theta)\big|_{\theta=\tau(\eta_{1}^{\alpha})}\cdot\operatorname{Vol}(\Theta^{\alpha})$ (S36)
where we used $\prod_{i\in I_{\alpha}}\partial_{i}\tau_{i}(\eta_{(i)})=\prod_{i\in I_{\alpha}}b_{i}-a_{i}=\operatorname{Vol}(\Theta^{\alpha})$. Denote by $h(\theta)$ the mixed partial derivative $\partial^{\alpha}g(\theta)$ to see that $h(\tau(\eta_{1}^{\alpha}))=\partial^{\alpha}g(\theta)|_{\theta=\tau(\eta_{1}^{\alpha})}$. Combining (S35) and (S36) gives that
$V^{\dagger}(g\circ\tau)=\sum_{\alpha\neq 0}\int_{[0,1]^{\alpha}}\operatorname{Vol}(\Theta^{\alpha})\left|h(\tau(\eta_{1}^{\alpha}))\right|d\eta^{\alpha}.$ (S37)
@@
$V^{\dagger}(g\circ\tau)=\sum_{\alpha\neq 0}\int_{\Theta^{\alpha}}\operatorname{Vol}(\Theta^{\alpha})\left|h(\tau(\eta_{1}^{\alpha}))\right|d\eta^{\alpha}=\sum_{\alpha\neq 0}\int_{\Theta^{\alpha}}\left|\partial^{\alpha}g(\theta_{*}^{\alpha})\right|d\theta^{\alpha}=V(g).$ (S38)
where $\operatorname{Vol}(\Theta^{\alpha})$ is canceled by the change of variables. Therefore, the inequality (S34) is rewritten as
$(*)\leq V(g)\operatorname{D}^{\dagger}(\{\eta_{n}\}_{n=1}^{K})$ (S39)
@@
$(*_{1})$ $=\sup_{\|f\|_{\mathcal{H}_{\kappa}\leq 1}}\left|\int_{\Theta}f(\theta)q(\theta)d\theta-\frac{\operatorname{Vol}(\Theta)}{N+M}\sum_{n=1}^{N+M}f(\theta_{n})q(\theta_{n})\right|\leq\operatorname{Vol}(\Theta)\underbrace{\sup_{\|f\|_{\mathcal{H}_{\kappa}\leq 1}}V(f\cdot q)}_{=:(*_{11})}\operatorname{D}^{*}(\{\theta_{n}\}_{n=1}^{M+N})$ (S41)
@@
$V(f\cdot q)$ $\leq\sum_{\alpha\neq 0}\operatorname{Vol}(\Theta^{\alpha})\sup_{\theta\in\Theta^{\alpha}}\left|\partial^{\alpha}(f(\theta_{*}^{\alpha})q(\theta_{*}^{\alpha}))\right|\leq\sum_{\alpha\neq 0}\operatorname{Vol}(\Theta^{\alpha})\underbrace{\sup_{\theta\in\Theta}\left|\partial^{\alpha}(f(\theta)q(\theta))\right|}_{=:(*_{12})}.$ (S42)
@@
$\partial^{\alpha}(f(\theta)q(\theta))=\sum_{\beta\leq\alpha}(\partial^{\alpha-\beta}f(\theta))(\partial^{\beta}q(\theta))$ (S43)
@@
$(*_{12})\leq\sup_{\theta\in\Theta}\sum_{\beta\leq\alpha}\left|(\partial^{\alpha-\beta}f(\theta))(\partial^{\beta}q(\theta))\right|\leq\sum_{\beta\leq\alpha}C_{q}^{\beta}.$ (S44)
@@
$(*_{11})\leq\sum_{\alpha\neq 0}\operatorname{Vol}(\Theta^{\alpha})\sum_{\beta\leq\alpha}C_{q}^{\beta}=:A_{1}$ (S45)
@@
$(*_{2})=\left|\int_{\Theta}q(\theta)d\theta-\frac{\operatorname{Vol}(\Theta)}{N+M}\sum_{n=1}^{N+M}q(\theta_{n})\right|\leq\operatorname{Vol}(\Theta)V(q)\operatorname{D}^{*}(\{\theta_{n}\}_{n=1}^{M+N})$ (S46)
@@
$V(q)$ $\leq\sum_{\alpha\neq 0}\operatorname{Vol}(\Theta^{\alpha})\sup_{\theta\in\Theta^{\alpha}}\left|\partial^{\alpha}q(\theta_{*}^{\alpha})\right|\leq\sum_{\alpha\neq 0}\operatorname{Vol}(\Theta^{\alpha})\sup_{\theta\in\Theta}\left|\partial^{\alpha}q(\theta)\right|.$ (S47)
@@
$V(q)$ $\leq\sum_{\alpha\neq 0}\operatorname{Vol}(\Theta^{\alpha})C_{q}^{\alpha}=:A_{2}$ (S48)
@@
$(*)$ $\leq\operatorname{Vol}(\Theta)A_{1}B_{1}\operatorname{D}^{*}(\{\theta_{n}\}_{n=1}^{M+N})+\operatorname{Vol}(\Theta)A_{2}B_{2}\operatorname{D}^{*}(\{\theta_{n}\}_{n=1}^{M+N})+B_{3}\frac{M}{N+M}.$ (S49)
@@ B.2 Proof of Corollary˜1
Recall that $\operatorname{D}^{\dagger}$ is the star discrepancy for the unit cube $0,1]^{d}$ defined in (S8) in Section˜A.2. The scaled Halton sequence $\{\theta_{i}\}_{i=1}^{\infty}$ is defined as an image of the original Halton sequence $\{\eta_{i}\}_{i=1}^{\infty}$ under the bijective affine map. By Lemma˜3, we have $\operatorname{D}^{*}\left(\{\theta_{n}\}_{n=1}^{K}\right)=\operatorname{D}^{\dagger}\left(\{\eta_{n}\}_{n=1}^{K}\right)$ for any number of points $K$. By Theorem 3.6 of Niederreiter1992, there exists some constant $C>0$ s.t. the star discrepancy $\operatorname{D}^{\dagger}$ of the original Halton sequence $\{\eta_{n}\}_{n=1}^{K}$ satisfies that
$\operatorname{D}^{\dagger}\left(\{\eta_{n}\}_{n=1}^{K}\right)\leq C\frac{\log(K)^{d}}{K}.$ (S50)
@@ B.3 Proof of Corollary˜2
By assumption, the kernel $\kappa$ is bounded, continuous, and integrally strictly positive definite. Hence, the resulting MMD metrizes the weak convergence by Theorem 7 of Simon-Gabriel2023. By Corollary˜1, the MMD between $\delta_{N}$ and $p$ converges to zero. The consistency of the Monte Carlo estimator immediately follows from the definition of the weak convergence. Here, since the domain $\Theta$ is compact, the continuous integrand $f$ is bounded by extreme value theorem. ∎
B.4 Proof of Theorem˜2
The proof uses Lemma˜4 presented in Section˜B.1, which holds for any proposal sequence.
@@ B.4 Proof of Theorem˜2
$I_{M+N}(f)$ $=\left|\int_{\Theta}f(\theta)g(\theta)u(\theta)d\theta-\frac{1}{N+M}\sum_{n=1}^{N+M}f(\theta_{n})g(\theta_{n})\right|.$ (S52)
@@
$(*_{1})$ $=\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}\left|\mathbb{E}_{\theta\sim u}\left[f(\theta)g(\theta)\right]-\mathbb{E}_{\theta\sim u_{M+N}}\left[f(\theta)g(\theta)\right]\right|=\sup_{h\in\mathcal{H}}\left|\mathbb{E}_{\theta\sim u}\left[h(\theta)\right]-\mathbb{E}_{\theta\sim u_{M+N}}\left[h(\theta)\right]\right|.$ (S53)
@@
$\mathcal{R}_{M+N}(\mathcal{H})$ $:=\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}}\mathbb{E}_{\epsilon_{1},\dots,\epsilon_{M+N}}\left[\sup_{h\in\mathcal{H}}\left|\frac{1}{M+N}\sum_{n=1}^{M+N}\epsilon_{n}h(\theta_{n})\right|\right]$ (S54)
$=\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}}\mathbb{E}_{\epsilon_{1},\dots,\epsilon_{M+N}}\left[\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}\left|\frac{1}{M+N}\sum_{n=1}^{M+N}\epsilon_{n}f(\theta_{n})g(\theta_{n})\right|\right].$ (S55)
@@
$\sup_{\theta\in\Theta}|h(\theta)|\leq\sup_{\theta\in\Theta}|f(\theta)|\sup_{\theta\in\Theta}|g(\theta)|\leq\sup_{\theta\in\Theta}|g(\theta)|=b.$ (S56)
It then follows from Theorem 4.10 of Wainwright2019 that
$\mathbb{P}\left((*_{1})\leq 2\mathcal{R}_{M+N}(\mathcal{H})+\delta\right)\geq 1-\exp\left(-\frac{(M+N)\delta^{2}}{2b^{2}}\right)$ (S57)
@@
$\mathcal{R}_{M+N}(\mathcal{H})$ $=\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}}\mathbb{E}_{\epsilon_{1},\dots,\epsilon_{M+N}}\left[\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}\left|\frac{1}{M+N}\sum_{n=1}^{M+N}\epsilon_{n}\langle f(\cdot),\kappa(\cdot,\theta_{n})\rangle_{\mathcal{H}_{\kappa}}g(\theta_{n})\right|\right]$ (S58)
$=\frac{1}{M+N}\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}}\mathbb{E}_{\epsilon_{1},\dots,\epsilon_{M+N}}\left[\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}\left|\left\langle f(\cdot),\sum_{n=1}^{M+N}\epsilon_{n}\kappa(\cdot,\theta_{n})g(\theta_{n})\right\rangle_{\mathcal{H}_{\kappa}}\right|\right].$ (S59)
@@
$\mathcal{R}_{M+N}(\mathcal{H})$ $=\frac{1}{M+N}\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}}\mathbb{E}_{\epsilon_{1},\dots,\epsilon_{M+N}}\left[\left\|\sum_{n=1}^{M+N}\epsilon_{n}\kappa(\cdot,\theta_{n})g(\theta_{n})\right\|_{\mathcal{H}_{\kappa}}\right]$ (S60)
$\leq\frac{1}{M+N}\sqrt{\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}}\mathbb{E}_{\epsilon_{1},\dots,\epsilon_{M+N}}\left[\left\|\sum_{n=1}^{M+N}\epsilon_{n}\kappa(\cdot,\theta_{n})g(\theta_{n})\right\|_{\mathcal{H}_{\kappa}}^{2}\right]}$ (S61)
$=\frac{1}{M+N}\sqrt{\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}}\mathbb{E}_{\epsilon_{1},\dots,\epsilon_{M+N}}\left[\sum_{n=1}^{M+N}\sum_{m=1}^{M+N}\epsilon_{n}\epsilon_{m}g(\theta_{n})\kappa(\theta_{n},\theta_{m})g(\theta_{m})\right]}.$ (S62)
@@
$\mathcal{R}_{M+N}(\mathcal{H})$ $\leq\frac{1}{M+N}\sqrt{\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}}\left[\sum_{n=1}^{M+N}g(\theta_{n})\kappa(\theta_{n},\theta_{n})g(\theta_{n})\right]}$ (S63)
$\leq\frac{1}{M+N}\sqrt{(M+N)b^{2}}\leq\frac{b}{\sqrt{M+N}}$ (S64)
where the second inequality follows from that $\kappa$ is bounded by $1$ and that $g(\theta)\leq\sup_{\theta\in\Theta}|g(\theta)|=b$. Plugging this upper bound of $\mathcal{R}_{M+N}(\mathcal{H})$ in the concentration inequality (S57) gives that
$\mathbb{P}\left((*_{1})\leq\frac{2b}{\sqrt{M+N}}+\delta\right)\geq 1-\exp\left(-\frac{(M+N)\delta^{2}}{2b^{2}}\right)\geq 1-2\exp\left(-\frac{(M+N)\delta^{2}}{2b^{2}}\right)$ (S65)
@@
$\mathbb{P}\left((*_{1})\leq\frac{2b+\sqrt{2\log(2/\epsilon)}b}{\sqrt{M+N}}\right)\geq 1-\epsilon$ (S66)
@@
$(*_{2})$ $=\left|\mathbb{E}_{\theta\sim u}\left[g(\theta)\right]-\mathbb{E}_{\theta\sim u_{M+N}}\left[g(\theta)\right]\right|=\left|\mathbb{E}_{\theta_{1},\dots,\theta_{M+N}\overset{i.i.d.}{\sim}u}[F(\theta_{1},\dots,\theta_{M+N})]-F(\theta_{1},\dots,\theta_{M+N})\right|.$ (S67)
Our aim is to apply the bounded difference inequality in Corollary 2.21 of Wainwright2019. The function $F$ admits the ‘bounded difference’ with respect to each $n$-th argument $\theta_{n}$. To see this, fix all arguments $\theta_{1},\dots,\theta_{M+N}$, except the $n$-th argument $\theta_{n}$, and observe for any $n$ that
$\left|F(\theta_{1},\dots,\theta_{n},\dots,\theta_{M+N})-F(\theta_{1},\dots,\theta_{n}^{\prime},\dots,\theta_{M+N})\right|=\frac{1}{M+N}\left|g(\theta_{n})-g(\theta_{n}^{\prime})\right|\leq\frac{2b}{M+N}$ (S68)
where $\theta_{n}$ and $\theta_{n}^{\prime}$ are arbitrary. Thus we can apply Corollary 2.21 of Wainwright2019 to see that
$\mathbb{P}\left((*_{2})\leq\delta\right)\geq 1-2\exp\left(-\frac{(M+N)\delta^{2}}{2b^{2}}\right)$ (S69)
@@
$\mathbb{P}\left((*_{2})\leq\frac{\sqrt{2\log(2/\epsilon)}b}{\sqrt{M+N}}\right)\geq 1-\epsilon.$ (S70)
@@
$\mathbb{P}\left(d(p,\delta_{N})\leq B_{1}\frac{2b+\sqrt{2\log(2/\epsilon)}b}{\sqrt{M+N}}+B_{2}\frac{\sqrt{2\log(2/\epsilon)}b}{\sqrt{M+N}}+B_{3}\frac{M}{N+M}\right)\geq 1-\epsilon$ (S71)